# EasyRec: Simple yet Effective Language Models for Recommendation

### v3

@@ 3.1.2. User Profiling
In practical scenarios, privacy concerns often limit the feasibility of generating user profiles based on demographic information. Instead, we can profile users by considering their collaborative relationships, using the generated profile information from their interacted items. This approach allows the user profiles to effectively capture the collaborative signals that reflect their preferences. Formally, the user profile generation process is:
(5) $\mathcal{P}_{u}=\text{LLM}(\mathcal{M}_{u},\{h_{i},\mathcal{P}_{i},r_{u,i}|i\in\mathcal{N}_{u}\}).$
Here, $\mathcal{M}_{u}$ represents the instruction for generating the user profile using a large language model (LLM). We sample a set of interacted items $\mathcal{N}_{u}$ from the user’s purchase history. We then combine the user’s feedback $r_{u,i}$ with the pre-generated item profiles $\mathcal{P}_{i}$ to create the user’s text description $\mathcal{P}_{u}$, which captures their preferences.
3.1.3. Advantages of Collaborative Profiling
@@
Our collaborative user and item profiling framework offers two key advantages for real-world recommendation scenarios which are elaborated as:
@@
  • Collaborative Information Preserved. Our collaborative profiling approach goes beyond just the original textual content, also capturing the semantics of user/item characteristics and their interaction patterns. By encoding these rich profiles into a shared feature space using a recommendation-oriented language model, the resulting embeddings of interacted users and items are brought closer together. This enables the recommenders to better identify relevant matches, even for ”zero-shot” users and items (those without prior interactions) which are ubiquitous in real-world scenarios. The system can leverage the collaborative signals encoded within the text-based profiles to make better recommendations, bridging the gap for these cold-start cases.
@@
  • Fast Adaptation to Dynamic Scenarios. Our user and item profiling approach, powered by robust language models, enables the recommender system to effectively handle the time-evolving nature of user preferences and interaction patterns. The key advantage is that simple updates to the textual user profiles can seamlessly reflect shifts in user interests and behaviors. This flexibility and responsiveness makes our approach well-suited for deploying recommender systems in dynamic environments where user interests can evolve over time.
3.2. Profile Embedder with Collaborative LM
@@
So far, we have generated rich textual profiles for users and items, moving beyond conventional ID-based embeddings. However, directly encoding these textual profiles into latent embeddings for making recommendations may have two key limitations:
@@
  • Capturing Recommendation-Specific Semantics. The text-based embeddings, while expressive, may not be optimized for the specific semantics and relationships most relevant to the recommendation task. For example, consider two item profiles: (1) “This user is passionate about advanced AI techniques, focusing on deep learning and AI research works”. (2) “With a passion for advanced AI development, this user delves into science fiction and AI-themed novels”. Though the profiles share textual similarity about AI, their target audiences differ - the first caters to AI scientists, the second to sci-fi readers. Directly encoding these profiles may overlook nuanced, recommendation-specific semantics. Refinement is needed to better align embeddings with the specific context and requirements of the recommendation system, beyond just textual similarity.
@@
  • Overlooking High-Order Collaborative Signals. While textual profiles offer rich semantic information, relying solely on them may cause us to overlook valuable high-order collaborative patterns that emerge from complex user-item interactions (Wang et al., 2019; Xia et al., 2023). These higher-order signals, such as transitive associations and community-level preferences, can provide additional insights that complement the textual data for user preference learning in recommender systems.
@@
To address these limitations, we propose a collaborative language modeling paradigm that seamlessly integrates the strengths of the semantic richness of the profiles and the valuable collaborative signals encoded from complex user-item interaction behaviors.
3.2.1. Bidirectional Transformer Encoder as Embedder
@@
We leverage a multi-layer bidirectional Transformer encoder as the embedder backbone, considering two key benefits: 1) Efficient Encoding: The encoder-only architecture focuses solely on generating effective text representations, enabling faster inference in recommendation systems. 2) Flexible Adaptation: By building on pre-trained Transformer models, we can leverage transfer learning to optimize the embedder for specific recommendation tasks.
@@
Let’s consider a user’s profile as a passage of $n$ words: $\mathcal{P}={w_{1},\ldots,w_{n}}$. We start by adding a special token [CLS] at the beginning of the word sequence. The tokenization layer then encodes the input sequence into initial embeddings, which serve as the input for the Transformer layers:
(6) $\{\mathbf{x}_{\texttt{[CLS]}}^{(0)},\ldots\mathbf{x}_{n}^{(0)}\}=\text{Tokenization}(\{w_{\texttt{[CLS]}},\ldots,w_{n}\}).$
@@
Here, $x^{(0)}\in\mathbb{R}^{d}$ is the embedding retrieved from the embedding table corresponding to the tokens, and the $(0)$ superscript indicates that it is the input to the $(0)$-th layer of the language model. The tokenization process also adds positional embeddings to the initial embeddings. The language model then encodes a sequence of final embeddings (one for each token):
(7) $\{\mathbf{e}_{[CLS]},\ldots,\mathbf{e}_{n}\}=\text{Encoder}(\{\mathbf{x}_{\texttt{[CLS]}}^{(0)},\ldots\mathbf{x}_{n}^{(0)}\}),$
@@
where Encoder($\cdot$) refers to the Transformer-based encoder-only language model. The key operation in the encoding process is the self-attention mechanism:
(8) $\text{Attention}(Q,K,V)=\text{softmax}({QK^{T}}/{\sqrt{d}})V$
[…]
@@
Here, $X\in\mathbb{R}^{n\times d}$ represents the stack of token embeddings, and $W^{Q/K/V}$ are the parameter matrices that map these embeddings into queries, keys, and values. This self-attention mechanism allows each token to aggregate information from all other tokens, ensuring that each token is informed about the entire sequence. Finally, we select the first embedding $\mathbf{e}_{\texttt{[CLS]}}$, which corresponds to the [CLS] token, as the representative embedding for the entire profile. This embedding is then passed through a multi-layer perceptron to obtain the final encoded representation $e$, as mentioned in Eq.(1):
(10) $\mathbf{e}=\text{MLP}(\mathbf{e}_{\texttt{[CLS]}})=\text{LM}(\mathcal{P}).$
@@
With these encoded text embeddings $\mathbf{e}$ for each user and item, we can predict the likelihood of interaction using cosine similarity and make recommendations as described in Eq.(2).
@@
Figure 3. Contrastive tuning of collaborative language model empowers the model to learn rich representations that closely align the text-based semantic space with the globally collaborative signals from user behavior patterns.
3.2.2. Collaborative LM with Contrastive Learning
@@
The motivation behind fine-tuning the collaborative language model using contrastive learning is to effectively capture and incorporate high-order collaborative signals into the recommendation model. Traditional recommenders using Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) optimize encoded embeddings with only one negative item per training sample. This approach limits the model’s ability to capture complex global user-item relationships.
@@
The supervised contrastive loss provides a strong alternative to traditional methods. By treating interacted user-item pairs as positive views and non-interacted pairs as negatives, it brings related item embeddings closer together in the feature space. As shown in Figure 3, contrastive learning incorporates a batch of negatives, allowing for a comprehensive adjustment of the encoded feature space, which enables the model to capture high-order collaborative relationships. To evaluate this, we conduct experiments in Section 4.2.3 to assess the impact of different learning objectives. Specifically, the supervised contrastive loss for tuning the language model can be expressed as follows:
(11) $\mathcal{L}_{\text{con}}=-\frac{1}{|\mathcal{B}|}\sum_{(u,i_{\text{pos}},i_{\text{neg}})}\log\frac{\exp(\cos(\mathbf{e}_{u},\mathbf{e}_{i_{\text{pos}}})/\tau)}{\sum_{m\in\mathcal{N}_{\text{neg}}}\exp(\cos(\mathbf{e}_{u},\mathbf{e}_{m})/\tau)},$
@@
Here, $\tau$ is a temperature hyperparameter that regulates the degree of contrastive learning, while $\mathcal{N}_{\text{neg}}$ denotes the set of in-batch negative items, which includes the negative item $i_{\text{neg}}$ and other items in the batch, excluding the positive item $i_{\text{pos}}$, with $|\mathcal{B}|$ depicts the batch size. Additionally, we follow prior work (Li et al., 2023; Hou et al., 2024) and incorporate an masked language modeling (MLM) loss $\mathcal{L}_{\text{mlm}}$. This technique randomly masks input tokens and trains the model to predict them, which helps stabilize the training process and enhance the model’s generalization ability. The final training objective is the combination of the supervised contrastive and the MLM loss:
(12) $\mathcal{L}=\mathcal{L}_{\text{con}}+\lambda\mathcal{L}_{\text{mlm}}.$
@@
where $\lambda$ is a hyperparameter that controls the balance between the contrastive learning and the masked language modeling loss.
3.3. Augmentation with Profile Diversification
The goal of our profile diversification approach is to enhance the model’s ability to generalize to unseen users and items. Representing each user or item with a single profile inherently limits the diversity of the representations, which can negatively impact the model’s performance and generalization. To address this, we propose augmenting the existing user/item profiles to allow for multiple profiles per entity. These augmented profiles capture the same semantic meaning, such as the personalized interaction preferences of users or the varied characteristics of items. Our two specific augmentation methods introduce controlled variations in the profiles while preserving the core semantic meaning.
@@ 3.3. Augmentation with Profile Diversification
Inspired by self-instruction mechanisms (Wang et al., 2023; Xu et al., 2024), large language models (LLMs) can be leveraged to rephrase user or item profiles while preserving their underlying meaning. This allows generating multiple semantically similar yet distinctly worded profiles from a single input. Applying this iterative rephrasing process can create a diverse set of augmented profiles, substantially expanding the available training data. This data augmentation technique is particularly valuable when the original dataset is limited, as the LLM-generated profiles can improve model generalization and robustness.
By leveraging large language models for profile diversification, we can create a set of diverse profiles for each user $u$ and item $i$.
(13) $\{\mathcal{P}\}_{u}$ =$=\{\mathcal{P}_{u};\,\,\mathcal{P}_{u}^{1},\,\mathcal{P}_{u}^{2},\,\ldots,\,\mathcal{P}_{u}^{t}\},$
(14) $\{\mathcal{P}\}_{i}$ =$=\{\mathcal{P}_{i};\,\,\mathcal{P}_{i}^{1},\,\mathcal{P}_{i}^{2},\,\ldots,\,\mathcal{P}_{i}^{t}\}.$
@@
Here, $\mathcal{P}_{u/i}$ represents the original profile, while $\mathcal{P}_{u/i}^{1-t}$ denotes the LLM-rephrased profiles, with $t$ indicating the number of diversification steps. During training, we randomly select a profile from the profile set for the user or item in each batch of user-item pairs.
4. Evaluation
This section evaluates the performance of the proposed EasyRec framework in addressing the following research questions (RQs):
@@ 4. Evaluation
  • RQ1: How effectively does the proposed EasyRec perform in matching unseen users and items (zero-shot) within text-based recommendation scenarios?
  • RQ2: How effectively does EasyRec integrate with and enhance various recommenders within text-based collaborative filtering scenarios, leveraging its capabilities as a language embedder?
  • RQ3: How effective is our proposed profile diversification mechanism for data augmentation in improving the performance of the recommendation language model?
  • RQ4: How well can our proposed text-based EasyRec paradigm adapt to accommodate changes in users’ dynamic preferences?
4.1. Experimental Settings
@@
Table 1. Statistics of the datasets, with ”Avg. n” representing the average number of interactions per user. Datasets marked with underline are from different platforms.
Datasets #Users #Items #Inters. Avg. n Density
Train Data 124,732 67,455 802,869 6.44 -
[…]
Test Data 55,877 28,988 615,210 11.01 -
[…]
@@
Table 2. The specifics of EasyRec’s model variants differ based on their model architecture and parameter sizes.
Model Layers Hidden size Heads Params
@@
Table 3. Text-based recommendation performance of various language models across different datasets. The best performance is indicated in bold, while the second-best is highlighted with underline. The script * denotes significance (p ¡ 0.05).
@@
Table 4. Comparison of EasyRec training with different optimization objectives (where ”Contrast” stands for contrastive).
Objective Sports Yelp
Recall@10 NDCG@10 Recall@10 NDCG@10
BPR Loss 0.0381 0.0226 0.0028 0.0021
Contrast Loss 0.0395 0.0236 0.0034 0.0026
4.1.1. Datasets
To assess our proposed model’s capability in encoding user/item textual profiles into embeddings for recommendation, we curated diverse datasets across various domains and platforms. A portion was used for training, while the remainder served as test sets for zero-shot evaluation. The dataset statistics are shown in Table 1. Due to the page limit, we place the detail of data resources are described in Appendix A.2.1
4.1.2. Evaluation Protocols
We employ two commonly used ranking-based evaluation metrics, Recall@$N$ and $N$, to assess performance in both text-based recommendation and collaborative filtering scenarios. Specifically, we compute these metrics for $N$ values of 10 and 20 (He et al., 2017, 2020). The evaluation is conducted using the all-rank protocol (He et al., 2020) with the predicted preference scores. In the context of text-based recommendation, particularly for datasets containing multiple LLM-diversified profiles (as discussed in Section 3.3), we calculate the metrics separately $t$ times based on different profile pairs $(\mathcal{P}_{u}^{1,\ldots,t},\mathcal{P}_{i}^{1,\ldots,t})$. Subsequently, we compute the mean value for each metric to obtain a comprehensive assessment. For the training datasets, we utilize the validation split for evaluation, while for the test datasets, we employ the test split.
4.2. Performance Comparision for Text-based Recommendation (RQ1)

### v4

@@ 3.1.2 User Profiling
In practical scenarios, privacy concerns limit generating user profiles from demographic information. Instead, we profile users by leveraging their collaborative relationships through interacted item profiles. This approach captures collaborative signals that reflect user preferences. The user profile generation process is defined as:
$\mathcal{P}_{u}=\text{LLM}(\mathcal{M}_{u},\{h_{i},\mathcal{P}_{i},r_{u,i}\mid i\in\mathcal{N}_{u}\}).$ (5)
Here, $\mathcal{M}_{u}$ is the instruction for using a LLM to generate the user profile. We sample interacted items $\mathcal{N}_{u}$ from the user’s purchase history. We combine their feedback $r_{u,i}$ with the pre-generated item profiles $\mathcal{P}_{i}$. This creates the user’s text description $\mathcal{P}_{u}$, capturing their preferences. The incorporation of sampled data ensures that LLMs accurately infer profiles. This includes item metadata and user reviews. The profiles authentically reflect users’ interaction preferences. To enhance profile quality, we adopt the principles of Chain of Thought (CoT)Wei et al. (2022) and Self-ConsistencyWang et al. (2023). We require LLMs to generate explanatory rationales alongside the profiles. It improves inference reliability. It also establishes an interpretable connection between collaborative signals and profile generation.
3.1.3 Advantages of Collaborative Profiling
@@
Our collaborative profiling framework offers two key advantages for real-world recommendation:
@@
  • Preservation of Collaborative Information. Our approach captures the original textual content. It also captures the semantics of user/item characteristics and their interaction patterns. We encode these profiles into a shared feature space. This uses a recommendation-oriented language model. The embeddings of interacted users and items are aligned. This allows recommenders to better identify relevant matches.
@@
  • Rapid Adaptation to Dynamic Scenarios. Our profiling enables the recommender system to adapt to evolving user preferences. It also adapts to changing interaction patterns. With robust language models, simple updates to textual user profiles can quickly reflect shifts in interests. They can also reflect shifts in behaviors. This flexibility makes our approach ideal for environments. User interests change over time in environments.
3.2 Profile Embedder with Collaborative LM
@@
We have generated rich textual profiles for users and items, moving beyond conventional ID-based embeddings. However, directly encoding these textual profiles into latent embeddings for recommendations has two key limitations:
@@
  • Capturing Recommendation-Specific Semantics. While text embeddings are expressive, they may not optimize for the specific semantics relevant to recommendations. For example, consider two user profiles: (i) "This user is passionate about advanced AI techniques, focusing on deep learning and research." (ii) "With a passion for advanced AI development, this user enjoys science fiction and AI-themed novels." Although both profiles mention AI, their target audiences differ—one caters to AI scientists, while the other targets sci-fi readers. Directly encoding these profiles may overlook recommendation-specific semantics, necessitating refinement to align embeddings with the context of recommendation.
@@
  • Overlooking High-Order Collaborative Signals. While textual profiles provide rich semantic information, relying solely on them may cause us to miss valuable high-order collaborative patterns. These patterns arise from complex user-item interactions Wang et al. (2019); Xia et al. (2023). Such signals include transitive associations and community-level preferences. They offer additional insights that enhance preference learning.
@@
To address these limitations, we propose a collaborative language modeling paradigm. It integrates the semantic richness of profiles with valuable collaborative signals from complex interactions.
3.2.1 Bidirectional Transformer Embedder
@@
We use a multi-layer bidirectional Transformer encoder as the backbone for two key benefits: 1) Efficient Encoding: The encoder-only architecture generates effective text representations, allowing faster inference in recommendation. 2) Flexible Adaptation: Leveraging pre-trained Transformer models enables us to optimize the embedder for specific recommendation tasks effectively.
@@
Let’s consider a user’s profile as a passage of $n$ words: $\mathcal{P}={w_{1},\ldots,w_{n}}$. We start by adding a special token [CLS] at the beginning of the word sequence. The tokenization layer $\text{Tok}(\cdot)$ then encodes the input sequence into initial embeddings, which serve as the input for the Transformer layers:
$\{\mathbf{x}_{\texttt{[CLS]}}^{(0)},\ldots\mathbf{x}_{n}^{(0)}\}=\text{Tok}(\{w_{\texttt{[CLS]}},\ldots,w_{n}\}).$ (6)
@@
Here, $x^{(0)}\in\mathbb{R}^{d}$ is the embedding from the embedding table for the tokens, with the $(0)$ superscript indicating it is the input to the $(0)$-th layer of the language model. The tokenization process also adds positional embeddings. The language model then encodes a sequence of final embeddings:
$\{\mathbf{e}_{[CLS]},\ldots,\mathbf{e}_{n}\}=\text{Enc}(\{\mathbf{x}_{\texttt{[CLS]}}^{(0)},\ldots\mathbf{x}_{n}^{(0)}\}),$ (7)
@@
where $\text{Enc}(\cdot)$ refers to the Transformer-based encoder-only LM. The key operation in the encoding process is the self-attention mechanism:
$\begin{split}\text{Attention}(Q,K,V)=\text{softmax}({QK^{T}}/{\sqrt{d}})V\\ \textit{w.r.t.}\,Q=XW^{Q},\,K=XW^{K},\,V=XW^{V}.\end{split}$ (8)
@@
Here, $X\in\mathbb{R}^{n\times d}$ represents the stack of token embeddings, while $W^{Q/K/V}$ are parameter matrices that map these embeddings into queries, keys, and values. This self-attention mechanism enables each token to aggregate information from all others, ensuring awareness of the entire sequence. We then select the first embedding $\mathbf{e}_{\texttt{[CLS]}}$ as the representative embedding for the profile. This embedding is passed through a multi-layer perceptron to obtain the encoded representation $e$, as shown in Eq.(1):
$\mathbf{e}=\text{MLP}(\mathbf{e}_{\texttt{[CLS]}})=\text{LM}(\mathcal{P}).$ (9)
@@
With these encoded text embeddings $\mathbf{e}$ for each user and item, we can predict the score of interaction using cosine similarity and make recommendations as described in Eq.(2).
@@
Figure 3: Contrastive tuning of the collaborative LM enables it to learn rich representations. It aligns the text-based semantic space with global collaborative signals.
3.2.2 Contrastive Collaborative LM Learning
@@
We fine-tune the collaborative LM using contrastive learning to effectively capture high-order collaborative signals. Traditional recommenders using Bayesian Personalized Ranking (BPR) Rendle et al. (2012) optimize embeddings with only one negative item per sample. This limits their ability to capture complex global user-item relationships.
@@
The supervised contrastive loss offers a strong alternative to traditional methods. By treating interacted user-item pairs as positives and non-interacted pairs as negatives, it brings related item embeddings closer in the feature space. As shown in Figure 3, contrastive learning uses a batch of negatives for a comprehensive adjustment of the encoded space, enabling the model to capture high-order collaborative relationships. We evaluate this in Appendix E to assess the impact of different learning objectives. The learning objective for tuning the language model is expressed as follows:
$\mathcal{L}_{\text{con}}=-\sum\log\frac{\exp(s_{u,i^{+}}/\tau)}{\sum_{j\in\mathcal{N}^{-}}\exp(s_{u,j}/\tau)},$ (10)
@@
where $s_{u,j}=\cos(\mathbf{e}_{u},\mathbf{e}_{j})$ denotes the cosine similarity between user $u$ and item $j$, $i^{+}$ represents the positive item that user $u$ has interacted with, $\tau$ is a temperature hyperparameter that controls the degree of learning, and $\mathcal{N}^{-}$ is the set of in-batch negative items. We also build on prior work Li et al. (2023); Hou et al. (2024a) by incorporating an auxiliary masked language modeling (MLM) loss $\mathcal{L}_{\text{mlm}}$. This technique randomly masks input tokens, training the model to predict them, which stabilizes training and enhances generalization. The final training objective combines the contrastive loss and the MLM loss:
$\mathcal{L}=\mathcal{L}_{\text{con}}+\lambda\mathcal{L}_{\text{mlm}}.$ (11)
@@
$\lambda$ is a hyperparameter that balances contrastive learning with masked language modeling loss.
3.3 Augmentation with Profile Diversification
To enhance generalization to unseen users/items, we propose profile diversification. Single profiles per user or item limit representation diversity and training quality, hurting performance. Our augmentation creates multiple profiles per entity while preserving semantic meaning—capturing personalized preferences for users and varied characteristics.
@@ 3.3 Augmentation with Profile Diversification
Inspired by self-instruction mechanisms Wang et al. (2023); Xu et al. (2024), LLMs can rephrase user or item profiles while preserving their underlying meaning. This generates multiple semantically similar yet distinctly worded profiles from single inputs. Iterative rephrasing creates diverse augmented profiles, substantially expanding available training data. This technique proves particularly valuable for limited datasets, as LLM-generated profiles improve model generalization and robustness. Through LLM-based diversification, we create diverse profile sets for each user and item.
$\{\mathcal{P}\}_{u}$ $=\{\mathcal{P}_{u};\,\,\mathcal{P}_{u}^{1},\,\mathcal{P}_{u}^{2},\,\ldots,\,\mathcal{P}_{u}^{t}\},$ (12)
$\{\mathcal{P}\}_{i}$ $=\{\mathcal{P}_{i};\,\,\mathcal{P}_{i}^{1},\,\mathcal{P}_{i}^{2},\,\ldots,\,\mathcal{P}_{i}^{t}\}.$ (13)
@@
Here, $\mathcal{P}_{u/i}$ represents the original profile, while $\mathcal{P}_{u/i}^{1-t}$ denotes the LLM-rephrased profiles, with $t$ indicating the number of diversification steps. During training, we randomly select one profile from the user’s or item’s profile set for each batch data.
4 Evaluation
@@
We evaluate the EasyRec in addressing the following research questions (RQs): RQ1: How effectively does the EasyRec perform in matching unseen users and items (zero-shot) within text-based recommendation? RQ2: How effectively does EasyRec integrate to enhance recommendations in text-based collaborative filtering scenarios? RQ3: How effective is our profile diversification mechanism for augmenting data and improving the recommendation model’s performance? RQ4: How well can our proposed text-based EasyRec adapt to changes in users’ dynamic preferences?
4.1 Experimental Settings
@@
Table 1: Dataset statisics: "Avg. n" is the average interactions per user, and "Inters." stands for interactions. Datasets with underlines are from different platforms.
Datasets #Users #Items #Inters. Avg. n
Train Data 124,732 67,455 802,869 6.44
-Arts 14,470 8,537 96,328 6.66
-Games 17,397 8,330 120,255 6.91
-Movies 16,994 9,370 134,649 7.92
-Home 22,893 13,070 131,556 5.75
-Electronics 26,837 14,033 165,628 6.17
-Tools 26,141 14,155 154,453 5.91
Test Data 55,877 28,988 615,210 11.01
-Sports 21,476 12,741 132,400 6.17
-Steam 23,310 5,237 316,190 13.56
-Yelp 11,091 11,010 166,620 15.02
@@
Table 2: Model variants of EasyRec differ by parameter size. "HS" stands for hidden size.
Model Layers HS Heads Params
@@
Table 3: Text-based recommendation performance of various LMs across different datasets. The best performance is indicated in bold, while the second-best is highlighted with underline. The script * denotes significance (p < 0.05).
@@
4.1.1 Datasets
To assess our proposed model’s capability in encoding user/item textual profiles into embeddings for recommendation, we curated diverse datasets across various domains and platforms. A portion was used for training, while the remainder served as test sets for zero-shot evaluation. The dataset statistics are shown in Table 1. Due to the page limit, we place the detail of data resources are described in Appendix F.1
4.1.2 Evaluation Protocols
We employ two ranking-based evaluation metrics, Recall@$N$ and NDCG@$N$, to assess performance in both text-based recommendation and collaborative filtering scenarios. Specifically, we compute these metrics for $N$ values of 10 and 20 He et al. (2017, 2020). The evaluation is conducted using the all-rank protocol He et al. (2020) with the predicted preference scores. In the context of text-based recommendation, particularly for datasets containing multiple LLM-diversified profiles (as discussed in Section 3.3), we calculate the metrics separately $t$ times based on different profile pairs $(\mathcal{P}_{u}^{1,\ldots,t},\mathcal{P}_{i}^{1,\ldots,t})$. Subsequently, we compute the mean value for each metric to obtain a comprehensive assessment. For the training datasets, we utilize the validation split for evaluation, while for the test datasets, we employ the test split.
4.2 Performance Comparision for Text-based Recommendation (RQ1)